### v2

# Hausdorff vs Gromov–Hausdorff distances

Abstract.
Let $M$ be a closed Riemannian manifold and let $X\subseteq M$. If the sample $X$ is sufficiently dense relative to the curvature of $M$, then the Gromov–Hausdorff distance between $X$ and $M$ is bounded from below by half their Hausdorff distance, namely $d_{\mathrm{GH}}(X,M)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)$. The constant $\frac{1}{2}$ can be improved depending on the dimension and curvature of the manifold $M$, and obtains the optimal value $1$ in the case of the unit circle, meaning that if $X\subseteq S^{1}$ satisfies $d_{\mathrm{GH}}(X,S^{1})<\tfrac{\pi}{6}$, then $d_{\mathrm{GH}}(X,S^{1})=d_{\mathrm{H}}(X,S^{1})$. We also provide versions lower bounding the Gromov–Hausdorff distance $d_{\mathrm{GH}}(X,Y)$ between two subsets $X,Y\subseteq M$. Our proofs convert discontinuous functions between metric spaces into simplicial maps between Čech or Vietoris–Rips complexes. We then produce topological obstructions to the existence of certain maps using the nerve lemma and the fundamental class of the manifold, thus lower bounding the Gromov–Hausdorff distance.
Key words and phrases:
Gromov–Hausdorff distance, Hausdorff distance, Nerve lemma, Čech and Vietoris–Rips complexes
1. Introduction
The Gromov–Hausdorff distance, denoted $d_{\mathrm{GH}}(X,Y)$, gives rise to a natural dissimilarity measure between two abstract metric spaces $X$ and $Y$ not necessarily sitting in a common ambient space [23, 27, 28]. Despite its increasing popularity and usefulness in the theoretical and applied communities, much is yet to be discovered about the Gromov–Hausdorff distance.
One such question, for $X$ a subset of a metric space $M$, is how $d_{\mathrm{GH}}(X,M)$ relates to the Hausdorff distance $d_{\mathrm{H}}(X,M)$. This paper furthers the understanding of the relationship between the Hausdorff distance and Gromov–Hausdorff distance when $M$ is a closed Riemannian manifold. Furthermore, we provide new lower bounds on $d_{\mathrm{GH}}(X,Y)$ when $X$ and $Y$ are dense-enough subsets of a manifold $M$.
Figure 1. A manifold $M$, a subset $X$ (orange diamonds), and a subset $Y$ (blue circles).
By definition, the Gromov–Hausdorff distance between $X$ and $M$ is bounded from above by the Hausdorff distance between them: $d_{\mathrm{GH}}(X,M)\leq d_{\mathrm{H}}(X,M)$. We show that if the sample $X$ is sufficiently dense relative to the curvature of a manifold $M$, then the Gromov–Hausdorff distance is also bounded from below by half the Hausdorff distance: $d_{\mathrm{GH}}(X,M)\geq\frac{1}{2}\cdot d_{\mathrm{H}}(X,M)$. Thus, the asymptotics of the Gromov–Hausdorff distance, whose definition involves isometric embeddings into arbitrary metric spaces, are controlled by the Hausdorff distance in a fixed space. Furthermore, we prove that if $X$ and $Y$ are two subsets of $M$ and dense enough relative to the convexity radius of $M$, then $d_{\mathrm{GH}}(X,Y)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M)$ (see Figure 1).
Theorem 1.
Let $M$ be a connected, closed Riemannian manifold with convexity radius $\rho(M)$. Then for any $X,Y\subseteq M$, we have
  1. (a)
    $d_{\mathrm{GH}}(X,M)\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M),\tfrac{1}{6}\rho(M)\right\}$, and
  2. (b)
    $d_{\mathrm{GH}}(X,Y)\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\tfrac{1}{6}\rho(M)-\frac{2}{3}d_{\mathrm{H}}(Y,M)\right\}$.
The interpretation of part (a) in Theorem 1 is as follows. If $X\subseteq M$ is sufficiently dense (meaning $d_{\mathrm{GH}}(X,M)<\tfrac{1}{6}\rho(M)$), then $d_{\mathrm{GH}}(X,M)\geq\tfrac{1}{2}d_{\mathrm{H}}(X,M)$.
Part (b) above is trivially true unless $Y$ is more than twice as dense as $X$, i.e. unless $d_{\mathrm{H}}(Y,M)<\frac{1}{2}d_{\mathrm{H}}(X,M)$.
Though part (a) is obtained from (b) by letting $Y=M$, we first give a separate proof of (a) since it has a short proof (see Section 3) that is approachable even for newcomers to this research area. Our proof uses Čech simplicial complexes in order to produce topological obstructions to the existence of functions between $X$ and $M$ of low distortion and codistortion, thus lower bounding the Gromov–Hausdorff distance. The topology of Čech complexes appears in three main ways. First, it allows us to turn possibly discontinuous functions between $X$ and $M$ into continuous (simplicial) maps between Čech complexes on $X$ and on $M$. Second, our quantitative bound on $X\subseteq M$ being “sufficiently dense” depends on the convexity radius of the manifold $M$, which arises since the nerve lemma implies that the Čech complex of $M$ is homotopy equivalent to $M$ for scale parameters up to the convexity radius. Third, the nerve lemma [8, 12] implies that the Čech complex of $X\subseteq M$ does not recover the fundamental class of $M$ until the scale is at least the Hausdorff distance $d_{\mathrm{H}}(X,M)$. Our proof technique for Theorem 1(a) is flexible: the remaining results in our paper show that various terms in part (a) can be improved by incorporating fancier tools into the same proof strategy.
The constants in Theorem 1 are not optimal, and we improve them by incorporating Vietoris–Rips complexes into the proofs. Before showing how to improve the constants for a general manifold, we first show how to improve them in the case when $M=S^{1}$ is the circle.
Theorem 2.
For any subsets $X,Y\subseteq S^{1}$ of the circle of circumference $2\pi$, we have
  1. (a)
    $d_{\mathrm{GH}}(X,S^{1})\geq\min\left\{d_{\mathrm{H}}(X,S^{1}),\tfrac{\pi}{6}\right\}$, and
  2. (b)
    $d_{\mathrm{GH}}(X,Y)\geq\min\left\{d_{\mathrm{H}}(X,S^{1})-d_{\mathrm{H}}(Y,S^{1}),\tfrac{\pi}{6}-\tfrac{1}{2}d_{\mathrm{H}}(Y,S^{1})\right\}$.
Part (a) in Theorem 2 is obtained from part (b) simply by letting $Y=S^{1}$. Since the Gromov–Hausdorff distance is never larger than the Hausdorff distance, if $X\subseteq S^{1}$ is sufficiently dense (meaning $d_{\mathrm{GH}}(X,S^{1})<\tfrac{\pi}{6}$), then we have an equality $d_{\mathrm{GH}}(X,S^{1})=d_{\mathrm{H}}(X,S^{1})$ between the Gromov–Hausdorff distance and the Hausdorff distance.
When applying part (b) above, one should always choose $X$ and $Y$ with $d_{\mathrm{H}}(Y,S^{1})\leq d_{\mathrm{H}}(X,S^{1})$. If both $X,Y\subseteq S^{1}$ are sufficiently dense, i.e. $d_{\mathrm{H}}(Y,S^{1})\leq d_{\mathrm{H}}(X,S^{1})\leq\frac{\pi}{6}$, then the first term achieves the minimum, implying $d_{\mathrm{GH}}(X,Y)\geq|d_{\mathrm{H}}(X,S^{1})-d_{\mathrm{H}}(Y,S^{1})|$.
We next show how to improve the bounds in Theorem 1 for a general manifold $M$, not only for a circle. The first term to improve is the constant in front of the convexity radius $\rho(M)$, which we improve using Gromov’s filling radius [29, 40].
Theorem 3.
Let $M$ be a connected, closed Riemannian manifold with filling radius $\mathrm{FillRad}(M)$. Then for any $X,Y\subseteq M$, we have
  1. (a)
    $d_{\mathrm{GH}}(X,M)\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M),\tfrac{1}{2}\rho(M),\tfrac{1}{3}\mathrm{FillRad}(M)\right\}$, and
  2. (b)
    $d_{\mathrm{GH}}(X,Y)\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\tfrac{1}{2}\rho(M)-d_{\mathrm{H}}(Y,M),\tfrac{1}{3}\mathrm{FillRad}(M)-\tfrac{2}{3}d_{\mathrm{H}}(Y,M)\right\}$.
Part (a) in Theorem 3 is obtained from part (b) simply by letting $Y=M$.
Next, in Theorem 4 we show how to improve the factor of $\frac{1}{2}$ in front of the $d_{\mathrm{H}}(X,M)$ term in Theorem 1 to a constant $\alpha(n,\kappa)$ which depends only on the dimension $n$ of the manifold and on the upper bound $\kappa$ of the sectional curvatures on $M$. These constants are equal to $\alpha(n,\kappa)=\sqrt{\tfrac{n+1}{2n}}$ if $\kappa\leq 0$. The proof of Theorem 4 enables better bounds because it incorporates not only Čech complexes but also Vietoris–Rips complexes, allowing us to incorporate Jung’s theorem on manifolds [18, 20, 21, 22].
Theorem 4.
Let $M$ be a connected, closed Riemannian $n$-manifold with convexity radius $\rho(M)$ and an upper bound $\kappa$ on all sectional curvatures at all points of $M$. For any subsets $X,Y\subseteq M$,
$d_{\mathrm{GH}}(X,Y)\geq\min\left\{\alpha(n,\kappa)d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\frac{\alpha(n,\kappa)\tau-2d_{\mathrm{H}}(Y,M)}{2\alpha(n,\kappa)+2}\right\},$
where $\tau=\rho(M)$ if $\kappa\leq 0$ and $\tau=\min\left\{\rho(M),\tfrac{\pi}{2\sqrt{\kappa+1}}\right\}$ if $\kappa>0$.
These lower bounds on the Gromov–Hausdorff distance are quite different than the general setting, since in Theorem 5 we show that if $Z$ is a general compact metric space (not necessarily a manifold) and $X$ is a subset thereof (not necessarily sufficiently dense), then the ratio between $d_{\mathrm{GH}}(X,Z)$ and $d_{\mathrm{H}}(X,Z)$ can be arbitrarily far away from one.
Theorem 5.
For any $\varepsilon>0$, there exists a compact metric space $Z$ and a subset $X\subseteq Z$ with
$d_{\mathrm{GH}}(X,Z)<\varepsilon\cdot d_{\mathrm{H}}(X,Z).$
In general the Gromov–Hausdorff distance is NP-hard to compute or approximate [7, 46]. Though our paper is not algorithmic, it points to situations where one may be able to approximate Gromov–Hausdorff distances with the simpler Hausdorff distances.
We begin in Section 2 with background on Hausdorff distances, Gromov–Hausdorff distances, correspondences, Čech complexes, the nerve lemma, the convexity radius of a manifold, and the fundamental class of a manifold. In Section 3 we prove Theorem 1(a), namely that $d_{\mathrm{GH}}(X,M)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)$ for $X$ a sufficiently dense subset of a manifold $M$. We introduce Vietoris–Rips complexes, which we use to show the constant $\frac{1}{2}$ can be improved to $1$ in the case of the circle (Theorem 2), in Section 4. Section 5 is dedicated to the proof of Theorem 1(b) for two subsets $X,Y\subseteq M$. In Section 6 we define the filling radius of a manifold to prove Theorem 3 with a relaxed density condition. In Section 7 we provide background on Hausmann’s Theorem and Jung’s Theorem, which we use to prove our improved bound for a general manifold (Theorem 4) with constant $\alpha(n,\kappa)$ in front of the $d_{\mathrm{H}}(X,M)$ term. We show in Section 8 that for general metric space subsets the ratio between the Hausdorff and Gromov–Hausdorff distances can be arbitrarily far away from one (Theorem 5). In Section 9 we conclude with a discussion of how this project began, and a list of open questions that we hope are of interest to others.
2. Background and notation
We present here the basic background material and notation used throughout the paper. See the following textbooks for more details: [14] for metric spaces and the Hausdorff and Gromov–Hausdorff distance; [37] for simplicial complexes and the nerve lemma; [33] for $\mathbb{Z}/2$ homology and cohomology; [10] for Riemannian manifolds and curvatures.
Metric spaces
Let $(X,d)$ be a metric space. For any $x\in X$, we let $B(x;r)=\{x^{\prime}\in X~{}|~{}d(x,x^{\prime})<r\}$ denote the metric ball of radius $r$ about $x$. If needed for the sake of clarity, we will sometimes write $B_{X}(x;r)$ instead of $B(x;r)$. For any $r\geq 0$ and $A\subseteq X$, we let $B(A;r)=\cup_{x\in A}B(x;r)$ denote the union of metric balls of radius $r$ centered at the points of $A$.
The diameter of a subset $A\subseteq X$ is the supremum of all distances between pairs of points of $A$. More formally, the diameter is defined as
$\mathrm{diam}(A)=\sup\{d(x,x^{\prime})\mid x,x^{\prime}\in A\}.$
The diameter of a compact set $A$ is always finite.
Hausdorff distance
Let $Z$ be a metric space. If $X$ and $Y$ are two submetric spaces of $Z$ then the Hausdorff distance between $X$ and $Y$ is
$d^{Z}_{\mathrm{H}}(X,Y)=\inf\{r\geq 0\mid X\subseteq B(Y;r)\text{ and }Y\subseteq B(X;r)\}.$
In other words, the Hausdorff distance finds the infimal real number such that if we thicken $Y$ by $r$ it contains $X$ and if we thicken $X$ by $r$ it contains $Y$; if no such $r$ exists then the Hausdorff distance is infinity.
Gromov–Hausdorff distance
The Gromov–Hausdorff distance $d_{\mathrm{GH}}(X,Y)$ between two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ is defined as the infimum, over all metric spaces $Z$ and isometric embeddings $\gamma\colon X\to Z$ and $\varphi\colon Y\to Z$, of the Hausdorff distance in $Z$ between $\gamma(X)$ and $\varphi(Y)$ [23, 27, 28, 48]. To see that one can avoid taking an infimum over a proper class, one can restrict attention to the case where $Z$ is the disjoint union of $X$ and $Y$, equipped with a metric extending the metrics on $X$ and $Y$. The Gromov–Hausdorff distance is a commonly used quantity in geometry and data analysis [5, 16, 17, 30, 31, 41, 45, 49]. Unlike the Hausdorff distance, the Gromov–Hausdorff distance considers sets $X$ and $Y$ that are not part of the same metric space. However, to compute it we need to embed $X$ and $Y$ in different common metric spaces $Z$, and then take the infimum of the Hausdorff distance over all $Z$ and isometric embeddings. Like the Hausdorff distance, it is possible for the Gromov–Hausdorff distance to be infinity. The Gromov–Hausdorff distance indeed gives a metric on the isometry classes of compact metric spaces.
Note that for any two metric spaces $X$ and $Y$ contained in some common metric space $Z$, we always have
$d_{\mathrm{GH}}(X,Y)\leq d_{\mathrm{H}}^{Z}(X,Y).$ (1)
Correspondences and distortion
A relation $R\subseteq X\times Y$ is a correspondence if the following two conditions hold:
  • For every $x\in X$, there exists some $y\in Y$ such that $(x,y)\in R$, and
  • For every $y\in Y$, there exists some $x\in X$ such that $(x,y)\in R$.
So every point in X is related to at least one point in Y, and vice-versa. If $C$ is a correspondence between $(X,d_{X})$ and $(Y,d_{Y})$, then the distortion of $C$ is:
$\mathrm{dis}(C)=\sup_{(x,y)\in C,(x^{\prime},y^{\prime})\in C}\left|d_{X}(x,x^{\prime})-d_{Y}(y,y^{\prime})\right|.$
We can equivalently define the Gromov–Hausdorff distance between two metric spaces $X$ and $Y$ to be
$d_{\mathrm{GH}}(X,Y)=\tfrac{1}{2}\inf_{C\subseteq X\times Y}\mathrm{dis}(C),$
where the infimum is taken over all correspondences between $X$ and $Y$ [14, 36].
If $X$ and $Y$ are both compact metric spaces, then the infimum in the definition of the Gromov–Hausdorf distance is attained [35].
Simplicial complexes.
For notational convenience, we identify simplicial complexes with their geometric realizations. A simplicial map $f\colon K\to L$ is continuous on geometric realizations. If $K$ and $L$ are simplicial complexes and $f,g\colon K\to L$ are simplicial maps, then $f$ and $g$ are said to be contiguous if for every simplex $\sigma\in K$, the union $f(\sigma)\cup g(\sigma)$ is contained in some simplex in $L$. Contiguous simplicial maps induce homotopic maps on their geometric realizations.
Figure 2. A finite set $X$ of points in the plane, along with three Čech complexes $\mathrm{\check{C}}\left(X;r\right)$ with increasing values of $r>0$.
For $Z$ a metric space, $X\subseteq Z$, and $r>0$, the Čech complex $\mathrm{\check{C}}\left(X;r\right)$ is defined as the simplicial complex having $X$ as its vertex set, and a finite subset $[x_{0},\ldots,x_{k}]\subseteq X$ as a simplex if there is some $z\in Z$ with $d(z,x_{i})<r$ for all $0\leq i\leq k$, i.e. if $\cap_{i=0}^{k}B_{Z}(x_{i};r)\neq\emptyset$. See Figure 2 for an example. We emphasize that this is the ambient Čech complex, since we are looking at intersections of balls in $Z$. The case when $Z=X$ is often referred to as the intrinsic Čech complex, but this case is not used in our paper.
Discontinuous functions induce simplicial maps
Čech complexes are related to the theory of Gromov–Hausdorff distances because they allow us to turn discontinuous functions between metric spaces into continuous maps between simplicial complexes. Indeed, we refer the reader to [15] for more information on the following lemma.
Lemma 2.1 (Correspondences and ambient Čech complexes [15]).
Let $Z$ be a metric space, let $X\subseteq Z$, and let $r>2\cdot d_{\mathrm{GH}}(X,Z)$. Then for any $\varepsilon>0$, there exist simplicial maps
[…]
such that the composition $\overline{g}\circ\overline{h}$ is contiguous to the inclusion $\iota:\mathrm{\check{C}}\left(Z;\varepsilon\right)\hookrightarrow\mathrm{\check{C}}\left(Z;3r+2\varepsilon\right)$, where all complexes are ambient Čech complexes using balls in $Z$.
Proof.
Since $2\cdot d_{\mathrm{GH}}(X,Z)<r$, there exists a correspondence $C\subseteq X\times Z$ such that $\mathrm{dis}(C)<r$. We can define (possibly discontinuous) functions $h\colon Z\to X$ and $g\colon X\to Z$ such that $(h(z),z)\in C$ for any $z\in Z$ and $(x,g(x))\in C$ for any $x\in X$.
We first show that the map $h$ extends to a simplicial map $\overline{h}\colon\mathrm{\check{C}}\left(Z;\varepsilon\right)\to\mathrm{\check{C}}\left(X;r+\varepsilon\right)$ on ambient Čech complexes defined via $\overline{h}([z_{0},\ldots,z_{k}])=[h(z_{0}),\ldots,h(z_{k})]$. Indeed, $[z_{0},\ldots,z_{k}]\in\mathrm{\check{C}}\left(Z;\varepsilon\right)$ implies that there exists some $z\in Z$ with $d_{Z}(z,z_{i})<\varepsilon$ for $0\leq i\leq k$. Since $(h(z),z),(h(z_{i}),z_{i})\in C$,
$d_{X}(h(z),h(z_{i}))\leq d_{Z}(z,z_{i})+r<r+\varepsilon.$
So, $d_{X}(h(z),h(z_{i}))<r+\varepsilon$ for all $i$, meaning $\overline{h}([z_{0},\dots,z_{k}])\in\mathrm{\check{C}}\left(X;r+\varepsilon\right)$. So $\overline{h}$ is a simplicial map.
We now show that $g$ also extends to a simplicial map $\overline{g}\colon\mathrm{\check{C}}\left(X;r+\varepsilon\right)\to\mathrm{\check{C}}\left(Z;3r+2\varepsilon\right)$. Note $[x_{0},\ldots,x_{k}]\in\mathrm{\check{C}}\left(X;r+\varepsilon\right)$ implies that there exists some $z\in Z$ (since we are using ambient Čech complexes in $Z$) with $d_{Z}(z,x_{i})<r+\varepsilon$ for $0\leq i\leq k$. From the triangle inequality, for any $0\leq i\leq k$ we get
$d_{Z}(x_{0},x_{i})\leq d_{Z}(x_{0},z)+d_{Z}(z,x_{i})<(r+\varepsilon)+(r+\varepsilon)=2r+2\varepsilon.$
Since $(x_{0},g(x_{0})),(x_{i},g(x_{i}))\in C$, we therefore get
$d_{Z}(g(x_{0}),g(x_{i}))\leq d_{X}(x_{0},x_{i})+r<(2r+2\varepsilon)+r=3r+2\varepsilon.$
So, $d_{Z}(g(x_{0}),g(x_{i}))<3r+2\varepsilon$ for all $i$, meaning $\overline{g}([x_{0},\dots,x_{k}])\in\mathrm{\check{C}}\left(Z;3r+2\varepsilon\right)$.
In order to show that the simplicial maps $\overline{g}\circ\overline{h}$ and $\iota$ are contiguous, let $\sigma=[z_{0},\ldots,z_{k}]\in\mathrm{\check{C}}\left(Z;\varepsilon\right)$. So there exists $z\in Z$ with $d_{Z}(z,z_{i})<\varepsilon$ for all $i$. Since $(h(z),z),(h(z_{i}),g(h(z_{i})))\in C$, by the definition of distortion we have
$d_{Z}(z,g(h(z_{i})))\leq d_{X}(h(z),h(z_{i}))+r\leq d_{Z}(z,z_{i})+2r<2r+\varepsilon\leq 3r+2\varepsilon.$
So $B_{Z}(z;3r+2\varepsilon)$ contains $z_{i}$ and $g(h(z_{i}))$ for all $i$, meaning $(\overline{g}\circ\overline{h})(\sigma)\cup\iota(\sigma)\in\mathrm{\check{C}}\left(Z;3r+2\varepsilon\right)$. ∎
Riemannian manifolds
Let $M$ be a Riemannian manifold without boundary. A subset $A\subseteq M$ is called geodesically convex if for any two points $p,q\in A$, there exists a unique minimizing geodesic segment from $p$ to $q$, and furthermore this geodesic is entirely in $A$.
Definition 2.2 (Convexity radius).
The convexity radius of $M$, denoted $\rho(M)$, is the supremum of the set of radii so that all smaller balls are geodesically convex. Formally,
$\rho(M)=\sup\;\{r\geq 0\mid B(p;s)\text{ is geodesically convex for all }0<s<r\text{ and }p\in M\}.$
If $M$ is compact, the convexity radius $\rho(M)$ is positive; see [10, Proposition 95]. For example, the unit sphere $S^{n}$ has convexity radius $\rho(S^{n})=\tfrac{\pi}{2}$.
The main theorems in this paper will be for a connected, closed Riemannian manifold $M$, where closed means compact and without boundary.
Nerve lemma
An open cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of a topological space $Z$ is a good cover if each set is contractible and all finite intersections of its elements are empty or contractible. The nerve of $\mathcal{U}$, denoted $\mathcal{N}(\mathcal{U})$, is defined to be the simplicial complex having $I$ as its vertex set, and a simplex $\{i_{0},\ldots,i_{k}\}$ when the finite intersection $U_{i_{0}}\cap\ldots\cap U_{i_{k}}$ is nonempty (see Figure 3). Under the right assumptions, the nerve preserves the homotopy type of the union $Z$, as stated by the following foundational result.
Figure 3. (Left) A good cover of a topological space and its nerve. (Right) A cover of a topological space that is not good, along with its nerve.
Lemma 2.3 (Nerve lemma [8, 12, 32]).
If $\mathcal{U}=\{U_{i}\}_{i\in I}$ is a good open cover of a paracompact topological space $Z$, then we have a homotopy equivalence $\mathcal{N}(\mathcal{U})\simeq Z$.
Furthermore, if $\mathcal{U}^{\prime}=\{U^{\prime}_{i}\}_{i\in I}$ is another good open cover of the paracompact space $Z$ with $U_{i}\subseteq U^{\prime}_{i}$ for all $i\in I$, then the natural inclusion $\iota\colon\mathcal{N}(\mathcal{U})\hookrightarrow\mathcal{N}(\mathcal{U}^{\prime})$ is a homotopy equivalence.
We refer to this latter statement as a functorial nerve lemma. It is implied by the standard proof of the nerve lemma as follows. When defining the map from $Z\to\mathcal{N}(\mathcal{U}^{\prime})$, use the same partition of unity subordinate to $\mathcal{U}$ used to obtain the map $Z\to\mathcal{N}(\mathcal{U})$. This produces the commutative diagram below, which implies that the natural inclusion $\iota$ is a homotopy equivalence. See for example [24, Lemma 3.7].
[…]
See [9] for further variants of the functorial nerve lemma, for example in the setting when the cover need not be good.
We note that if $X$ is a metric space and $A\subseteq X$, then the Čech complex is an example of a nerve complex, namely $\mathrm{\check{C}}\left(A;r\right)=\mathcal{N}(\{B_{X}(a;r)~{}|~{}a\in A\})$.
Corollary 2.4.
Let $M$ be a Riemannian manifold. Then $\mathrm{\check{C}}\left(M;r\right)\simeq M$ for all $r<\rho(M)$. Furthermore, if $0<r\leq r^{\prime}<\rho(M)$, then the natural inclusion $\iota\colon\mathrm{\check{C}}\left(M;r\right)\hookrightarrow\mathrm{\check{C}}\left(M;r^{\prime}\right)$ is a homotopy equivalence.
Proof.
Since $r<\rho(M)$, the ball $B(m;r)$ is geodesically convex for all $m\in M$. Furthermore, the open cover $\{B(m;r)\}_{m\in M}$ of $M$ is good, since an intersection of geodesically convex sets is geodesically convex, and since a nonempty geodesically convex set is contractible. And $M$ is paracompact since it is a metric space. Therefore, the Nerve Lemma (Lemma 2.3) implies $\mathrm{\check{C}}\left(M;r\right)\simeq M$.
Lemma 2.3 also implies that the inclusion map is a homotopy equivalence. ∎
We note the cover $\{B(m;r)\}_{m\in M}$ of $M$ is not locally finite. Nevertheless, since $M$ is paracompact, a (locally finite) partition of unity exists subordinate to this open cover. This idea is explained in [47, Theorem 13.1.3], and used in the proof of [32, Corollary 4G.3].
The fundamental class of a manifold
In this paper, we always take homology with $\mathbb{Z}/2$-coefficients and manifolds without boundary. Let $M$ be a connected (not necessarily compact) $n$–dimensional manifold without boundary. The $n$-dimensional homology $H_{n}(M;\mathbb{Z}/2)$ of $M$ is well-understood and is dictated by the compactness of $M$. More specifically,
$H_{n}(M;\mathbb{Z}/2)\simeq\begin{cases}\mathbb{Z}/2&\text{ if }M\text{ is compact}\\ 0&\text{ if }M\text{ is not compact}.\end{cases}$ (2)
To see this, we observe from the Poincare duality that
$H_{n}(M;\mathbb{Z}/2)\simeq H^{0}_{c}(M;\mathbb{Z}/2),$
where $H^{0}_{c}(M)$ denotes the 00-dimensional cohomology group with compact support. If $M$ is compact, cohomology with compact support matches the usual cohomology, and so $H^{0}_{c}(M,\mathbb{Z}/2)=\mathbb{Z}/2$. The fundamental class over $\mathbb{Z}/2$ of a compact, connected $n$–manifold $M$ is the (non-trivial) generator $[M]$ of the homology group $H_{n}(M,\mathbb{Z}/2)$; see [33]. In the case when $M$ is not compact, one can see that a 00-cocycle can not be supported on a compact subset of $M$, i.e., $H^{0}_{c}(M,\mathbb{Z}/2)=0$.
Lemma 2.5.
Let $M$ be a connected closed $n$-manifold. If $M^{\prime}$ is an open proper subset of $M$, then $H_{n}(M^{\prime})\cong 0$.
Proof.
The lemma is clearly true if $M^{\prime}$ is empty, so let $M^{\prime}\neq\emptyset$. The space $M^{\prime}$ is an $n$-manifold since it is an open subset of $M$. We further note that $M^{\prime}$ is not compact, since then it would be a nonempty proper subset that is both open and closed in $M$, contradicting the fact that $M$ is connected. The result then follows from (2). ∎
3. A bound in manifolds, with a non-optimal constant
We recall the statement of Theorem 1.
See 1
The purpose of this section is to give the proof of part (a). We defer the proof of part (b) until Section 5.
Proof of Theorem 1(a).
We will show that if $d_{\mathrm{GH}}(X,M)<\frac{1}{6}\rho(M)$, then $d_{\mathrm{GH}}(X,M)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)$.
Fix $r$ such that $2\cdot d_{\mathrm{GH}}(X,M)<r<\frac{1}{3}\rho(M)$, and fix $\varepsilon>0$ so that $3r+2\varepsilon<\rho(M)$. Every Čech complex in this proof is an ambient Čech complex with balls in $M$. By Lemma 2.1 we get simplicial maps
[…]
whose composition $\overline{g}\circ\overline{h}$ is homotopic to the inclusion $\iota\colon\mathrm{\check{C}}\left(M;\varepsilon\right)\hookrightarrow\mathrm{\check{C}}\left(M;3r+2\varepsilon\right)$. The two collections of balls $\{B(m;\varepsilon)\}_{m\in M}$ and $\{B(m;3r+2\varepsilon)\}_{m\in M}$ each form good open covers of $M$, since $0<\varepsilon<3r+2\varepsilon<\rho(M)$. The nerve lemma (Corollary 2.4) therefore implies that $\overline{g}\circ\overline{h}$ is a homotopy equivalence between spaces homotopy equivalent to $M$. Hence $\overline{g}_{*}\circ\overline{h}_{*}$ preserves the fundamental class, i.e. $\overline{g}_{*}\circ\overline{h}_{*}$ is an isomorphism on $n$-dimensional homology with $\mathbb{Z}/2$ coefficients, where $n$ is the dimension of the manifold $M$:
$\mathbb{Z}/2\cong H_{n}(\mathrm{\check{C}}\left(M;\varepsilon\right))\xrightarrow{\overline{h}_{*}}H_{n}(\mathrm{\check{C}}\left(X;r+\varepsilon\right))\xrightarrow{\overline{g}_{*}}H_{n}(\mathrm{\check{C}}\left(M;3r+2\varepsilon\right))\cong\mathbb{Z}/2.$
Figure 4. Picture of the balls determining $\mathrm{\check{C}}\left(X;r+\varepsilon\right)$ with $r+\varepsilon<d_{\mathrm{H}}(X,M)$ in the proof of Theorem 1(a); note that the balls do not cover $M$.
If we had $r+\varepsilon<d_{\mathrm{H}}(X,M)$, then the collection $\{B(x,r+\varepsilon)\}_{x\in X}$ of open balls would not cover $M$; see Figure 4. By the nerve lemma, $\mathrm{\check{C}}\left(X;r+\varepsilon\right)$ would be homotopy equivalent to an open proper subset of $M$, implying $H_{n}(\mathrm{\check{C}}\left(X;r+\varepsilon\right))=0$ by Lemma 2.5. This would contradict the fact that $(\overline{g}\circ\overline{h})_{*}$ is nonzero on $H_{n}$. Therefore, it must be that $r+\varepsilon\geq d_{\mathrm{H}}(X,M)$ for all $2\cdot d_{\mathrm{GH}}(X,M)<r<\frac{1}{3}\rho(M)$ and sufficiently small $\varepsilon>0$. Thus $2\cdot d_{\mathrm{GH}}(X,M)\geq d_{\mathrm{H}}(X,M)$, i.e. $d_{\mathrm{GH}}(X,M)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)$, as desired. ∎
4. Vietoris–Rips complexes and the case of the circle
We show how to improve the constant on the term $\frac{1}{2}d_{\mathrm{H}}(X,M)$ in the lower bounds in Theorem 1. We first do so in this section in the case when $M=S^{1}$ is the circle, when we improve the constant $\frac{1}{2}$ all the way to $1$, which cannot be improved further. Our main tool, Vietoris–Rips complexes, will also be important when improving the constant $\frac{1}{2}$ in the case of general manifolds $M$ in Section 7.
Vietoris–Rips complexes
For a metric space $(X,d)$ and scale $r>0$, the Vietoris–Rips complex $\mathrm{VR}\left(X;r\right)$ is defined as the simplicial complex having $X$ as its vertex set, and each finite subset $\sigma$ of $X$ with $\mathrm{diam}(\sigma)<r$ as a simplex. In other words, $[x_{0},\ldots,x_{k}]$ is a simplex if $d(x_{i},x_{j})<r$ for all $0\leq i,j\leq k$. See Figure 5.
Figure 5. A finite set $X$ of points in the plane, along with three Vietoris–Rips complexes $\mathrm{VR}\left(X;r\right)$ with increasing values of $r>0$.
Vietoris–Rips complexes are related to the theory of Gromov–Hausdorff distances because they allow us to turn discontinuous functions between metric spaces into continuous maps between simplicial complexes; this perspective from [15] was also recently used in [5].
Lemma 4.1 (Correspondences and Vietoris–Rips complexes [15]).
Let $X$ and $Y$ be metric spaces and let $r>2\cdot d_{\mathrm{GH}}(X,Y)$. Then for any $\varepsilon>0$, there exist simplicial maps:
[…]
such that the composition $\overline{h}\circ\overline{g}$ is contiguous to the inclusion $\mathrm{VR}\left(Y;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(Y;2r+\varepsilon\right)$.
We defer the proof to Appendix A, since it is analogous to the proof of Lemma 2.1. Later sections will also require the following lemma, which gives a tighter interleaving in the case of subset inclusion.
Lemma 4.2 (Hausdorff distance and Vietoris–Rips complexes).
Let $Z$ be a metric space, and let $Y\subseteq Z$ satisfy $r>2\cdot d_{\mathrm{H}}(Y,Z)$. Then for any $\varepsilon>0$, there exist simplicial maps:
[…]
such that the composition $\iota\circ\overline{f}$ is contiguous to the inclusion $\mathrm{VR}\left(Z;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(Z;r+\varepsilon\right)$.
We again defer the proof to Appendix A.
The case of the circle
Let $S^{1}$ be the circle with the geodesic metric and circumference $2\pi$. In Theorem 2, we use Vietoris–Rips complexes to improve the bound from Theorem 1 in the specific case of the circle. Indeed, we show how to improve the bound $d_{\mathrm{GH}}(X,S^{1})\geq\frac{1}{2}\cdot d_{\mathrm{H}}(X,S^{1})$ from Theorem 1 to be an equality, namely $d_{\mathrm{GH}}(X,S^{1})=d_{\mathrm{H}}(X,S^{1})$, for $X\subseteq S^{1}$ sufficiently dense.
First, we prove a lemma demonstrating the equivalence of Čech and Vietoris–Rips complexes at small scales.
Lemma 4.3 (Čech–Rips equivalence).
For any subset $X\subseteq S^{1}$ and $0<r<\frac{2\pi}{3}$, we have the simplicial inclusion $\mathrm{VR}\left(X;r\right)\hookrightarrow\mathrm{\check{C}}\left(X;\frac{r}{2}\right)$. As a consequence, $\mathrm{VR}\left(X;r\right)=\mathrm{\check{C}}\left(X;\frac{r}{2}\right)$.
Proof.
Let $\sigma=[x_{0},\ldots,x_{k}]\in\mathrm{VR}\left(X;r\right)$, which means $\mathrm{diam}(\sigma)<r$. Without any loss of generality, we assume that the geodesic distance between $x_{0}$ and $x_{k}$ equals $\mathrm{diam}(\sigma)$. Since $r<\frac{2\pi}{3}$, this implies that $x_{0}$ and $x_{k}$ are the unique points in $\sigma$ with $d(x_{0},x_{k})=\mathrm{diam}(\sigma)$, and furthermore that $\sigma$ is contained in the unique geodesic joining $x_{0}$ and $x_{k}$. If $y\in S^{1}$ is the midpoint of this geodesic arc, then the metric ball of radius $\frac{r}{2}$ centered at $y$ covers $\sigma$, giving $\sigma\in\mathrm{\check{C}}\left(X;\frac{r}{2}\right)$. This implies that $\mathrm{VR}\left(X;r\right)\hookrightarrow\mathrm{\check{C}}\left(X;\frac{r}{2}\right)$. It is always true that $\mathrm{\check{C}}\left(X;\frac{r}{2}\right)\hookrightarrow\mathrm{VR}\left(X;r\right)$, due to the triangle inequality, giving the other direction of the equality $\mathrm{VR}\left(X;r\right)=\mathrm{\check{C}}\left(X;\frac{r}{2}\right)$. ∎
This lemma will be generalized to the setting of a general manifold in Lemma 7.4, where we relate it to Jung’s theorem [18]. In the proof of Lemma 4.3, the fact that $\sigma$ is contained in the geodesic joining $x_{0}$ and $x_{k}$ reminds us of a local version of the definition of $\delta$-hyperbolicity [13] (small enough triangles in the circle are $\delta$-slim with $\delta=0$), and also of a local version of Helly’s theorem [18, 19] (small enough arcs in the circle that intersect pairwise have a point of common intersection).
See 2
Proof of Theorem 2.
Part (a) of Theorem 2 is obtained from part (b) by letting $Y=S^{1}$. Hence it suffices to prove (b).
Let $\delta=d_{\mathrm{H}}(Y,S^{1})$. We will show if $d_{\mathrm{GH}}(X,Y)<\frac{\pi}{6}-\frac{\delta}{2}$, then $d_{\mathrm{GH}}(X,Y)\geq d_{\mathrm{H}}(X,S^{1})-\delta$.
Fix $r$ such that $2\cdot d_{\mathrm{GH}}(X,Y)<r<\frac{\pi}{3}-\delta$, and fix $\varepsilon>0$ so that $r+\delta+\varepsilon<\frac{\pi}{3}$. By Lemma 4.1, we get continuous maps
[…]
whose composition $\overline{g}\circ\overline{h}$ is homotopic to the inclusion $\iota:\mathrm{VR}\left(Y;2\delta+2\varepsilon\right)\hookrightarrow\mathrm{VR}\left(Y;2r+2\delta+2\varepsilon\right)$. Since $2r+2\delta+2\varepsilon<\frac{2\pi}{3}$, using Lemma 4.3 we can write these complexes as
[…]
The two collections of balls $\{B(y;\delta+\varepsilon)\}_{y\in Y}$ and $\{B(y;r+\delta+\varepsilon)\}_{y\in Y}$ each form good open covers of $S^{1}$, since $d_{\mathrm{H}}(Y,M)<\delta+\varepsilon<r+\delta+\varepsilon<\frac{\pi}{3}<\frac{\pi}{2}=\rho(S^{1})$. The nerve lemma (Corollary 2.4) therefore implies that $\iota$ and $\overline{g}\circ\overline{h}$ are each homotopy equivalences between spaces homotopy equivalent to $S^{1}$. So $\overline{g}_{*}\circ\overline{h}_{*}$ is an isomorphism on $1$-dimensional homology.
If we had $\frac{r}{2}+\delta+\varepsilon<d_{\mathrm{H}}(X,S^{1})$, then the collection of open balls $\left\{B\left(x,\frac{r}{2}+\delta+\varepsilon\right)\right\}_{x\in X}$ would not cover $S^{1}$. By the nerve lemma, $\mathrm{\check{C}}\left(X;\frac{r}{2}+\delta+\varepsilon\right)$ would be homotopy equivalent to an open proper subset of $S^{1}$, implying $H_{1}\left(\mathrm{\check{C}}\left(X;\frac{r}{2}+\delta+\varepsilon\right)\right)=0$ by Lemma 2.5. This would contradict the fact that $(\overline{g}\circ\overline{h})_{*}$ is nonzero on $H_{1}(\bullet)$. Therefore, it must be that $\frac{r}{2}+\delta+\varepsilon\geq d_{\mathrm{H}}(X,S^{1})$ for all $2\cdot d_{\mathrm{GH}}(X,Y)<r<\frac{\pi}{3}-\delta$ and sufficiently small $\varepsilon>0$. So $d_{\mathrm{GH}}(X,Y)+\delta\geq d_{\mathrm{H}}(X,S^{1})$, giving $d_{\mathrm{GH}}(X,Y)\geq d_{\mathrm{H}}(X,S^{1})-d_{\mathrm{GH}}(Y,S^{1})$, as desired. ∎
Instead of passing to Čech complexes, the above proof could have instead used only Vietoris–Rips complexes, using the winding fraction [1, 3, 4] to determine when the Vietoris–Rips complex of a subset of the circle is homotopy equivalent to the circle or not, and the fact that $\mathrm{VR}\left(S^{1};r\right)\simeq S^{1}$ for $r<\frac{2\pi}{3}$. This is how our first proof of Theorem 2(a), the first result we proved during this research project, proceeded.
5. A bound for two subsets of a manifold using Hausmann’s theorem
Before producing a lower bound on the Gromov–Hausdorff distance between two subsets of a general manifold, we review Hausmann’s theorem. The basic idea is that Hausmann’s theorem is an analogue of the nerve lemma for Vietoris–Rips complexes instead of Čech complexes.
Hausmann’s theorem
The following result by Hausmann states that $M$ is homotopy equivalent to its Vietoris–Rips complex for a scale smaller than the convexity radius.
Theorem 5.1 (Hausmann’s Theorem [34]).
For any $0<r<\rho(M)$, we have a homotopy equivalence $\mathrm{VR}\left(M;r\right)\simeq M$.
Furthermore, if $0<r\leq r^{\prime}<\rho(M)$, then the inclusion $\mathrm{VR}\left(M;r\right)\hookrightarrow\mathrm{VR}\left(M;r^{\prime}\right)$ is a homotopy equivalence [50].
See [2] for an alternate proof of Hausmann’s theorem in the setting of Vietoris–Rips metric thickenings, which (with the $<$ convention) are now known to be weakly homotopy equivalent to the Vietoris–Rips simplicial complexes [26, 25, 6].
Two subsets of a manifold
We now prove Theorem 1(b) for two subsets of a manifold, which states the following. Let $M$ be a connected, closed Riemannian manifold with convexity radius $\rho(M)$. Then for any $X,Y\subseteq M$, we have
$d_{\mathrm{GH}}(X,Y)\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\tfrac{1}{6}\rho(M)-\tfrac{2}{3}d_{\mathrm{H}}(Y,M)\right\}.$
Proof of Theorem 1(b).
Let $0<\delta<2\cdot d_{\mathrm{H}}(Y,M)$. We will show if $d_{\mathrm{GH}}(X,Y)<\frac{1}{6}\rho(M)-\frac{2}{3}d_{\mathrm{H}}(Y,M)$, then $d_{\mathrm{GH}}(X,Y)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)-\frac{\delta}{2}$.
Note that
$2\cdot d_{\mathrm{GH}}(X,Y)<\tfrac{1}{3}\rho(M)-\tfrac{4}{3}d_{\mathrm{H}}(Y,M)<\tfrac{1}{3}\rho(M)-\tfrac{2}{3}\delta.$
So, we can fix $r$ such that $2\cdot d_{\mathrm{GH}}(X,Y)<r<\frac{1}{3}\rho(M)-\frac{2}{3}\delta$, and fix $\varepsilon>0$ so that $3r+2\delta+2\varepsilon<\rho(M)$. We produce the following sequence of continuous maps. The natural inclusion of the Vietoris–Rips complex into the Čech complex at the same scale gives the inclusion $\iota$. The inclusion $\iota^{\prime}$ always holds for Čech and Vietoris–Rips complexes after doubling the scale. We note that $\iota^{\prime}\circ\iota$ is equal to an inclusion (dashed), i.e., the middle triangle commutes. Since $2\cdot d_{\mathrm{GH}}(X,Y)<r$, we use Lemma 4.1 to get the maps $\overline{h}$ and $\overline{g}$, which satisfy that $\overline{g}\circ\iota^{\prime}\circ\iota\circ\overline{h}$ is homotopic to the inclusion map.
[…]
Finally, since $\delta<2\cdot d_{\mathrm{H}}(Y,M)$, Lemma 4.2 gives the maps $\overline{f}$ and $\iota^{\prime\prime}$, which satisfy that $\iota^{\prime\prime}\circ\overline{g}\circ\iota^{\prime}\circ\iota\circ\overline{h}\circ\overline{f}$ is homotopic to the inclusion $\mathrm{VR}\left(M;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(M;3r+2\delta+2\varepsilon\right)$. By Hausmann’s theorem (Theorem 5.1) and the fact that $\varepsilon<3r+2\delta+2\varepsilon<\rho(M)$, this inclusion is a homotopy equivalence between spaces homotopy equivalent to $M$. Thus the composition $\iota^{\prime\prime}\circ\overline{g}\circ\iota^{\prime}\circ\iota\circ\overline{h}\circ\overline{f}$ preserves the nonzero fundamental class in $n$-dimensional homology, where $n$ is the dimension of the manifold $M$.
If we had $r+\delta+\varepsilon<d_{\mathrm{H}}(X,M)$, then the collection $\left\{B\left(x,r+\delta+\varepsilon\right)\right\}_{x\in X}$ of open balls would not cover $M$. By the nerve lemma, $\mathrm{\check{C}}\left(X;r+\delta+\varepsilon\right)$ would be homotopy equivalent to an open proper subset of $M$, implying $H_{n}\left(\mathrm{\check{C}}\left(X;r+\delta+\varepsilon\right)\right)=0$ by Lemma 2.5, a contradiction. Therefore, it must be that $r+\delta+\varepsilon\geq d_{\mathrm{H}}(X,M)$ for all $2\cdot d_{\mathrm{GH}}(X,Y)<r<\frac{1}{3}\rho(M)-\frac{2}{3}\delta$ with $\varepsilon>0$ sufficiently small. So, $2\cdot d_{\mathrm{GH}}(X,Y)+\delta\geq d_{\mathrm{H}}(X,M)$, giving $d_{\mathrm{GH}}(X,Y)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)-\frac{\delta}{2}\geq\frac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M)$. ∎
Instead of considering Vietoris–Rips complexes of the manifold $M$ in the first and last steps of the sequence of maps in the proof of Theorem 1(b), it might be possible to instead apply Latschev’s theorem [38, 39, 42].
Figure 6. In both images, the subset $X$ (orange diamonds) consists of two points at distance $1$ apart, and same for the subset $Y$ (blue circles). (Left) A tree with four edges of length $\frac{1}{2}$ and one edge of length $1$. Note $d_{\mathrm{H}}(X,M)=2=4\cdot\frac{1}{2}=4\cdot d_{\mathrm{H}}(Y,M)$. (Right) Riemannian surface $M$ is the boundary of the an $\varepsilon$-thickening of the tree in $\mathbb{R}^{3}$, equipped with the geodesic metric. Note $d_{\mathrm{H}}(X,M)\approx 4\cdot d_{\mathrm{H}}(Y,M)$.
A particular case of Theorem 1(b) says the following. If $X$ and $Y$ are two isometric subsets of a connected, closed Riemannian manifold $M$, and if $Y$ is close enough to $M$ (meaning $d_{\mathrm{H}}(Y,M)<\frac{1}{4}\rho(M)$), then the Hausdorff distance from $X$ to $M$ and from $Y$ to $M$ cannot differ by more than a factor of two: $d_{\mathrm{H}}(X,M)\leq 2\cdot d_{\mathrm{H}}(Y,M)$. Indeed, if $X$ and $Y$ are isometric, then $d_{\mathrm{GH}}(X,Y)=0$, which can only be the case if the minimum on the right hand side of Theorem 1(b) is zero. And, $d_{\mathrm{H}}(Y,M)<\frac{1}{4}\rho(M)$ implies that the second term in this minimum is positive, meaning that the first term in the minimum must be nonnegative, yielding $d_{\mathrm{H}}(X,M)\leq 2\cdot d_{\mathrm{H}}(Y,M)$. If $X$ and $Y$ are singletons, then the triangle inequality immediately gives $d_{\mathrm{H}}(X,M)\leq 2\cdot d_{\mathrm{H}}(Y,M)$, without any need for the density assumption $d_{\mathrm{H}}(Y,M)<\frac{1}{4}\rho(M)$. However, once the isometric subsets $X$ and $Y$ consist of at least two points, the density assumption $d_{\mathrm{H}}(Y,M)<\frac{1}{4}\rho(M)$ is needed in general. Indeed, see the example in Figure 6(right), in which $X$ and $Y$ each consist of two points in $M$ at unit distance apart, in which $Y$ is not dense enough compared to the convexity radius of $M$, and in which $d_{\mathrm{H}}(X,M)\approx 4\cdot d_{\mathrm{H}}(Y,M)$. Of course, the full power of Theorem 1(b) is that it can provide lower bounds on $d_{\mathrm{GH}}(X,Y)$ even when $X$ and $Y$ are not isometric; Figure 6 simply shows the necessity of a density assumption.
6. Improving the convexity radius term via the filling radius
In [29], Gromov defines the filling radius of a manifold as follows. Any metric space $(Z,d_{Z})$ can be isometrically embedded into the space $L^{\infty}(Z)$ of all functions on $Z$, equipped with the supremum metric; this is called the Kuratowski embedding $\iota\colon Z\hookrightarrow L^{\infty}(Z)$ defined via $z\mapsto d_{Z}(z,\cdot)$. The filling radius of an $n$-dimensional manifold $M$ is the infimal scale parameter $r>0$ such that the inclusion $M\hookrightarrow B_{L^{\infty}(X)}(\iota(M);r)$ induces a non-injective homomorphism on $n$-dimensional homology, i.e., kills the fundamental class of $M$. Here $B_{L^{\infty}(X)}(M;r)$ is the union of all balls in $L^{\infty}(X)$ of radius $r$ that are centered at a point in $\iota(M)$. In [40], Lim, Mémoli, and Okutan prove that the filling radius, denoted $\mathrm{FillRad}(M)$, can equivalently be defined as the supremal scale parameter $r>0$ such that the inclusion $\mathrm{VR}\left(M;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(M;2r\right)$ induces an isomorphism on $n$-dimensional homology for all sufficiently small $\varepsilon>0$. See [40, Proposition 9.6] for an application to Gromov–Hausdorff distances.
Figure 7. A two-dimensional torus $M$ whose convexity radius is small (determined by the thinnest part of the neck) and whose filling radius is large.
Theorem 3 replaces the $\frac{1}{6}\rho(M)$ convexity radius terms from Theorem 1 with a larger $\frac{1}{2}\rho(M)$ term. The price we pay is introducing a new $\frac{1}{3}\mathrm{FillRad}(M)$ filling radius term. In general, the filling radius $\mathrm{FillRad}(M)$ can be much larger than the convexity radius $\rho(M)$; see Figure 7. That is why Theorem 3 can provide a large improvement over Theorem 1.
See 3
Proof.
Part (a) is obtained from part (b) by letting $Y=M$. Hence we prove (b). We follow the same notation and almost the same argument as the proof of Theorem 1(b). As before, let $0<\delta<2\cdot d_{\mathrm{H}}(Y,M)$, and assume that $d_{\mathrm{GH}}(X,Y)$ is strictly smaller than the last two terms in the minimum. Choose $r$ such that
$2\cdot d_{\mathrm{GH}}(X,Y)<r<\min\left\{\rho(M)-\delta,\tfrac{2}{3}\mathrm{FillRad}(M)-\tfrac{2}{3}\delta\right\},$
and choose $\varepsilon>0$ such that $r+\delta+\varepsilon<\rho(M)$ and $3r+2\delta+2\varepsilon<2\cdot\mathrm{FillRad}(M)$. We get the same diagram:
[…]
The composition is still homotopic to the inclusion $\mathrm{VR}\left(M;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(M;3r+2\delta+2\varepsilon\right)$, which by the bound on the filling radius and [40] induces an isomorphism on the fundamental class, i.e. on the $n$-dimensional homology group. Furthermore, note that $r+\delta+\varepsilon<\rho(M)$ implies that the Čech complex $\mathrm{\check{C}}\left(X;r+\delta+\varepsilon\right)$ is homotopy equivalent to the union $B(X,r+\delta+\varepsilon)$. Arguing as before, we obtain $d_{\mathrm{GH}}(X,Y)\geq\frac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M)$, as desired. ∎
7. Improving the Hausdorff distance constant via Jung’s theorem
We review Jung’s theorem, which will allow us to improve the constant on the term $\frac{1}{2}d_{\mathrm{H}}(X,M)$ in Theorem 1 to a better constant in Theorem 4. Jung’s theorem allows us to produce a tighter interleaving between Čech and Vietoris–Rips complexes.
Jung’s theorem
The definition of sectional curvatures of an abstract manifold $M$ uses a lot of machinery from Riemannian geometry. We skip the definition here, referring the reader to a textbook on the subject, e.g., [11, Chapter 9]. For a point $p\in M$ and unit norm vectors $u,v\in T_{p}(M)$ in the tangent space of $M$ at $p$, the sectional curvature at $p$ along the plane spanned by $u$ and $v$ is denoted by $\kappa_{p}(u,v)$. Intuitively, it measures the Gaussian curvature at $p$ for the $2$-dimensional submanifold with tangent plane spanned by $u$ and $v$. For example, the $n$-sphere of radius $R$ has a constant sectional curvature of $1/R^{2}$.
Let $\kappa(M)\in\mathbb{R}$ denote the supremum of the set of sectional curvatures $\kappa_{p}(u,v)$ across all $p\in M$ and all $u,v\in T_{p}(M)$. Since $M$ is compact, it can be shown that $\kappa(M)$ is bounded; see for example [11, p. 166].
Jung’s theorem for manifolds
For a compact subset $A\subseteq M$, the diameter satisfies $\mathrm{diam}(A)<\infty$. We define its circumradius in $M$ to be
$R(A)\coloneqq\inf_{m\in M}\max_{a\in A}d_{M}(a,m).$
Intuitively, the circumradius is the radius of a smallest closed metric ball (when one exists) in $M$ that contains $A$.
A point $c\in M$ satisfying $\max_{a\in A}d_{M}(a,c)=R(A)$ is called a circumcenter of $A$, and is denoted by $c(A)$. For $A$ compact, the circumradius is uniquely defined, but a circumcenter may not exist. When $M=\mathbb{R}^{n}$, however, the circumcenter exists uniquely. Moreover, for a compact Euclidean subset $A\subset\mathbb{R}^{d}$, the classical Jung’s theorem [18, Theorem 2.6] states that $R(A)\leq\sqrt{\tfrac{n}{2(n+1)}}~{}\mathrm{diam}(A)$.
Jung’s theorem was further extended by Dekster in [20, 21, 22], first for compact subsets of Riemannian manifolds with constant sectional curvatures, and then for Alexadrov spaces of curvature bounded above. A corollary in [22, Section 2] affirms that the circumcenter $c(A)$ exists (possibly non-uniquely) for a compact set $A$ if:
  1. (1)
    $A$ is contained in the interior of a compact geodesically convex domain $C\subseteq M$, and
  2. (2)
    $\mathrm{diam}(C)<\frac{2\pi}{3\sqrt{\kappa(M)}}$ if $\kappa(M)>0$.
Moreover, $c(A)$ belongs to the interior of $C$, and we have the following bound on the circumradius of $A$ in terms of the diameter of $A$.
Theorem 7.1 (Extented Jung’s theorem [22]).
Let $M$ be a compact, connected, $n$–dimensional manifold with sectional curvatures bounded above by $\kappa\in\mathbb{R}$. Let $A\subseteq M$ be compact with $\mathrm{diam}(A)<\rho(M)$ if $\kappa\leq 0$ and $\mathrm{diam}(A)<\min\left\{\rho(M),\tfrac{2\pi}{3\sqrt{\kappa}}\right\}$ if $\kappa>0$. Then the circumcenter $c(A)$ exists in $M$, and
$\mathrm{diam}(A)\geq\begin{cases}\frac{2}{\sqrt{-\kappa}}\sinh^{-1}\left(\sqrt{\frac{n+1}{2n}}\sinh\left(\sqrt{-\kappa}\;R(A)\right)\right)&\text{for }\kappa<0\\ 2R(A)\sqrt{\frac{n+1}{2n}}&\text{for }\kappa=0\\ \frac{2}{\sqrt{\kappa}}\sin^{-1}\left(\sqrt{\frac{n+1}{2n}}\sin\left(\sqrt{\kappa}\;R(A)\right)\right)&\text{for }\kappa>0\text{ and }R(A)\in\left[0,\tfrac{\pi}{2\sqrt{\kappa}}\right].\end{cases}$
This theorem (and the proof of Lemma 7.3 in Appendix B) leads us to define the following constants.
Definition 7.2.
For each integer $n\geq 1$ and $\kappa\in\mathbb{R}$, we define
$\alpha(n,\kappa)=\begin{cases}\sqrt{\tfrac{n+1}{2n}}&\text{for }\kappa\leq 0\\ \sqrt{\tfrac{n+1}{2n}}\sin\left(\tfrac{\pi}{2}\sqrt{\frac{\kappa}{\kappa+1}}\right)/\left(\tfrac{\pi}{2}\sqrt{\frac{\kappa}{\kappa+1}}\right)&\text{for }\kappa>0.\end{cases}$
Note that $\alpha(n,\kappa)$ is continuous as a function of $\kappa$, since $\alpha(n,\kappa)\to\sqrt{\frac{n+1}{2n}}$ as $\kappa\to 0$.
Note that $\sqrt{\tfrac{n+1}{2n}}$ is a decreasing function of $n$ that approaches its infimum value $\frac{1}{\sqrt{2}}$ as $n\to\infty$. So for $\kappa\leq 0$, we have $\frac{1}{\sqrt{2}}\leq\alpha(n,\kappa)\leq 1$. For $\kappa>0$, we note that the function $\sin\left(\tfrac{\pi}{2}\sqrt{\frac{\kappa}{\kappa+1}}\right)/\left(\tfrac{\pi}{2}\sqrt{\frac{\kappa}{\kappa+1}}\right)$ is a decreasing function of $\kappa$ that approaches its infimumum value $\frac{2}{\pi}$ as $\kappa\to\infty$. Therefore, for $n,\kappa$ arbitrary, we have $0.45\approx\frac{\sqrt{2}}{\pi}=\frac{2}{\pi}\cdot\frac{1}{\sqrt{2}}\leq\alpha(n,\kappa)\leq 1$. Our Theorem 4 only improves upon the $\frac{1}{2}\rho(M)$ term in Theorem 1 when $\alpha(n,\kappa)\geq\frac{1}{2}$, but this happens for a large range of $n$ and $\kappa$ values, and in particular whenever $\kappa\leq 0$.
Lemma 7.3 (Circumradius).
Let $A\subseteq M$ be compact with $\mathrm{diam}(A)<\tau$, where $\tau=\rho(M)$ for $\kappa\leq 0$ and $\tau=\min\left\{\rho(M),\tfrac{\pi}{2\sqrt{\kappa+1}}\right\}$ for $\kappa>0$. Then the circumcenter $c(A)$ exists in $M$, and the diameter and circumradius of $A$ satisfy $\mathrm{diam}(A)\geq 2\alpha(n,\kappa)\cdot R(A)$.
See Appendix B for the proof. The following lemma generalizes Lemma 4.3 from the circle case to general manifolds.
Figure 8. An inclusion of simplicial complexes $\mathrm{\check{C}}\left(X;\frac{r}{2}\right)\hookrightarrow\mathrm{VR}\left(X;r\right)$. We also have an inclusion $\mathrm{VR}\left(X;r\right)\hookrightarrow\mathrm{\check{C}}\left(X;\tfrac{r}{2\alpha(n,\kappa)}\right)$ obtained by increasing the size of the balls.
Lemma 7.4.
Let $M$ be a closed Riemannian manifold with $X\subseteq M$. For any $0<r<\tau$, we have the inclusion $\mathrm{VR}\left(X;r\right)\hookrightarrow\mathrm{\check{C}}\left(X;\tfrac{r}{2\alpha(n,\kappa)}\right)$, where $\tau=\rho(M)$ if $\kappa\leq 0$ and $\tau=\min\left\{\rho(M),\tfrac{\pi}{2\sqrt{\kappa+1}}\right\}$ if $\kappa>0$.
Proof.
Let $\sigma\in\mathrm{VR}\left(X;r\right)$, so $\mathrm{diam}(\sigma)<r<\tau$. From Lemma 7.3, it follows that $c(\sigma)$ exists in $M$, moreover $R(\sigma)\leq\frac{\mathrm{diam}(\sigma)}{2\alpha(n,\kappa)}<\frac{r}{2\alpha(n,\kappa)}$. Consequently, $\sigma$ must be contained in the metric ball of radius $\frac{r}{2\alpha(n,\kappa)}$ centered at $c(\sigma)$. Therefore, $\sigma\in\mathrm{\check{C}}\left(X;\tfrac{r}{2\alpha(n,\kappa)}\right)$. ∎
Manifolds with an improved constant
The following theorem improves upon the constant $\frac{1}{2}$ in the Hausdorff distance term $\tfrac{1}{2}d_{\mathrm{H}}(X,M)$ in Theorem 1.
See 4
We note that $\alpha(n,\kappa)=\alpha(1,0)=1$ when $M=S^{1}$ is the unit circle.
Proof.
Let $0<\delta<2\cdot d_{\mathrm{H}}(Y,M)$ and let $\alpha=\alpha(n,\kappa)$. We will show if $d_{\mathrm{GH}}(X,Y)<\frac{\alpha\tau-\delta}{2\alpha+2}$, then $d_{\mathrm{GH}}(X,Y)\geq\alpha\cdot d_{\mathrm{H}}(X,M)-\frac{\delta}{2}$.
Fix $r$ such that $2\cdot d_{\mathrm{GH}}(X,Y)<r<\frac{\alpha\tau-\delta}{\alpha+1}$, and fix $\varepsilon>0$ so that
$\frac{(\alpha+1)r+\delta+\varepsilon}{\alpha}<\tau.$ (3)
We produce the following sequence of continuous maps. Since $r+\delta+\varepsilon<(\alpha+1)r+\delta+\varepsilon<\alpha\tau\leq\tau$ (where the middle step is from (3)), Lemma 7.4 gives the inclusion $\iota$. After doubling the scale, the inclusion $\iota^{\prime}$ always holds for Čech and Vietoris–Rips complexes. We note that $\iota^{\prime}\circ\iota$ is equal to the dashed inclusion, so the middle triangle commutes. Since $2\cdot d_{\mathrm{GH}}(X,Y)<r$, we use Lemma 4.1 to get the maps $\overline{h}$ and $\overline{g}$, which satisfy that $\overline{g}\circ\iota^{\prime}\circ\iota\circ\overline{h}$ is homotopic to the inclusion map.
[…]
Finally, Lemma 4.2 gives the maps $\overline{f}$ and $\iota^{\prime\prime}$, which satisfy that $\iota^{\prime\prime}\circ\overline{g}\circ\iota^{\prime}\circ\iota\circ\overline{h}\circ\overline{f}$ is homotopic to the inclusion $\mathrm{VR}\left(M;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(M;\tfrac{(\alpha+1)r+\delta+\varepsilon}{\alpha}\right)$. By Hausmann’s theorem (Theorem 5.1) and the fact that $0<\varepsilon<\tfrac{(\alpha+1)r+\delta+\varepsilon}{\alpha}<\tau\leq\rho(M)$ from (3), this inclusion is a homotopy equivalence between spaces homotopy equivalent to $M$. Hence the composition $\iota^{\prime\prime}\circ\overline{g}\circ\iota^{\prime}\circ\iota\circ\overline{h}\circ\overline{f}$ preserves the fundamental class in $n$-dimensional homology.
If we had $\frac{r+\delta+\varepsilon}{2\alpha}<d_{\mathrm{H}}(X,M)$, then the collection $\left\{B\left(x,\frac{r+\delta+\varepsilon}{2\alpha}\right)\right\}_{x\in X}$ of open balls would not cover $M$. By the nerve lemma, $\mathrm{\check{C}}\left(X;\frac{r+\delta+\varepsilon}{2\alpha}\right)$ would be homotopy equivalent to an open proper subset of $M$, implying $H_{n}\left(\mathrm{\check{C}}\left(X;\frac{r+\delta+\varepsilon}{2\alpha}\right)\right)=0$ by Lemma 2.5, contradicting the fact that the fundamental class is preserved. Therefore, it must be that $\frac{r+\delta+\varepsilon}{2\alpha}\geq d_{\mathrm{H}}(X,M)$ for all $2\cdot d_{\mathrm{GH}}(X,Y)<r<\frac{\alpha\tau-\delta}{\alpha+1}$ and sufficiently small $\varepsilon>0$. So, $2\cdot d_{\mathrm{GH}}(X,Y)+\delta\geq 2\alpha\cdot d_{\mathrm{H}}(X,M)$. Thus, $d_{\mathrm{GH}}(X,Y)\geq\alpha\cdot d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M)$. ∎
Example 7.5.
The $n$-dimensional flat torus $T^{n}=S^{1}\times\ldots\times S^{1}$ is endowed with a Riemannian metric that makes the sectional curvatures at each point zero, giving $\alpha(n,\kappa)=\sqrt{\tfrac{n+1}{2n}}$. The convexity radius is $\rho(T^{n})=\frac{\pi}{2}$. Theorem 4 gives that for any $X,Y\subseteq T^{n}$,
$d_{\mathrm{GH}}(X,Y)\geq\min\left\{\sqrt{\tfrac{n+1}{2n}}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\frac{\sqrt{\tfrac{n+1}{2n}}\tfrac{\pi}{2}-2d_{\mathrm{H}}(Y,M)}{2\sqrt{\tfrac{n+1}{2n}}+2}\right\}.$
8. Ratio can be arbitrarily small in general
We next show that in general, the Hausdorff and Gromov–Hausdorff distances between general compact metric spaces and a subspace thereof can be arbitrarily far apart. Indeed, in Theorem 5 we show that for any $\varepsilon>0$, there exists a compact metric space $Z$ and a subset $X\subseteq Z$ with $d_{\mathrm{GH}}(X,Z)<\varepsilon\cdot d_{\mathrm{H}}(X,Z)$.
The “compact” hypothesis makes this result meaningful, since compact metric spaces do not admit isometries to proper subsets. By contrast, note that the non-compact metric space $[0,\infty)$ with the Euclidean metric has an isometric subspace $[n,\infty)$ satisfying $d_{\mathrm{GH}}([0,\infty),[n,\infty))=0$ and $d_{\mathrm{H}}([0,\infty),[n,\infty))=n$ for all $n>0$.
See 5
Proof.
We consider $\mathbb{R}^{n}$ with the Euclidean $\ell_{2}$-metric. Denote the standard basis vectors in $\mathbb{R}^{n}$ by $e_{1},\dots,e_{n}$, and let $x_{j}=\sum_{i=1}^{j}i\cdot e_{i}$  for $j\in\{1,\ldots,n\}$. Now let $Z=\{x_{1},\dots,x_{n}\}$ and $X=\{x_{1},\dots,x_{n-1}\}\subseteq Z$. For the Hausdorff distance between $X$ and $Z$, we have that $d_{\mathrm{H}}(X,Z)=\|x_{n}-x_{n-1}\|=n$. Consider the linear map $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined by $f(e_{i})=e_{i+1}$ for $i\in\{1,\dots,n-1\}$ and $f(e_{n})=e_{1}$. For $j\in\{2,\dots,n\}$, we compute $\|x_{j}-f(x_{j-1})\|=\|\sum_{i=1}^{j}e_{i}\|=\sqrt{j}$. Further, $\|x_{1}-f(x_{1})\|=\|e_{1}-e_{2}\|=\sqrt{2}$. Thus we get that $d_{\mathrm{H}}(f(X),Z)=\sqrt{n}$. Since $f$ is an isometry of $\mathbb{R}^{n}$ this proves that ${d_{\mathrm{GH}}(X,Z)\leq\sqrt{n}}$. This proves the theorem, by picking $n$ large enough so that $\frac{1}{\sqrt{n}}<\varepsilon$. ∎
Remark 8.1.
The proof above shows that the ratio between the Hausdorff distance of compact subsets $X\subseteq Z$ in $\mathbb{R}^{n}$ and the infimal Hausdorff distance of isometric copies of these subsets in $\mathbb{R}^{n}$ may be arbitrarily far from one. The sets $X$ and $Z$ in the proof above can be replaced by the compact simplices $\mathrm{conv}\ X$ and $\mathrm{conv}\ Z$ with the same Hausdorff distances, that is, $d_{\mathrm{H}}(\mathrm{conv}\ X,\mathrm{conv}\ Z)=n$, while $d_{\mathrm{H}}(f(\mathrm{conv}\ X),\mathrm{conv}\ Z)=\sqrt{n}$. We note that $\mathrm{conv}\ X$ and $\mathrm{conv}\ Z$ are manifolds with boundary.
9. Conclusion and open questions
We have furthered the study of the relationship between the Hausdorff and Gromov–Hausdorff distances, particularly for sufficiently dense subsets of a closed Riemannian manifold. We conclude with a discussion of how this project arose, along with a list of open questions.
In our paper, we have used the convenient definition of the Gromov–Hausdorff distance as $d_{\mathrm{GH}}(X,Y)=\tfrac{1}{2}\inf_{C\subseteq X\times Y}\mathrm{dis}(C)$, where the infimum is taken over all correspondences between $X$ and $Y$. It was observed in [36] that the Gromov–Hausdorff distance can be equivalently defined as
$d_{\mathrm{GH}}(X,Y)=\tfrac{1}{2}\inf_{g,h}\max\left\{\mathrm{dis}(g),\mathrm{dis}(h),\mathrm{codis}(g,h)\right\},$
where $g\colon X\to Y$ and $h\colon Y\to X$ are possibly discontinuous functions. Here the distortion of the function $g$ is
$\mathrm{dis}(g)=\sup_{x,x^{\prime}\in X}|d_{X}(x,x^{\prime})-d_{Y}(g(x),g(x^{\prime}))|$
(and similarly for $\mathrm{dis}(h)$), and the codistortion of the pair of functions $g$ and $h$ is
$\mathrm{codis}(g,h)=\sup_{x\in X,y\in Y}|d_{X}(x,h(y))-d_{Y}(g(x),y)|.$
The papers [41, 5] on Gromov–Hausdorff distances between spheres gave lower bounds on $d_{\mathrm{GH}}(X,Y)$ by lower bounding $\mathrm{dis}(h)$, the distortion of functions $Y\to X$, in particular in a setting where $Y$ was a higher-dimensional sphere, $X$ was a lower-dimensional sphere, and it sufficed to consider odd functions. Topological obstructions were used in both [41, 5] to bound $\mathrm{dis}(h)$, and [5] obtained these topological obstructions by first converting the (possibly discontinuous) function $h$ into a continuous map between Vietoris–Rips complexes. By contrast, our paper on Hausdorff versus Gromov–Hausdorff distances arose in part while looking for lower bounds on $d_{\mathrm{GH}}(X,Y)$ that also involved $\mathrm{codis}(g,h)$. We succeeded by noting that control on $\mathrm{codis}(g,h)$ allowed us to show that a (possibly discontinuous) composition $g\circ h\colon Y\to Y$ produced a continuous map on Vietoris–Rips complexes $\overline{g}\circ\overline{h}\colon\mathrm{VR}\left(Y;r\right)\to\mathrm{VR}\left(Y;r^{\prime}\right)$ with $r\leq r^{\prime}$ that was contiguous to the inclusion map $\mathrm{VR}\left(Y;r\right)\hookrightarrow\mathrm{VR}\left(Y;r^{\prime}\right)$ (see Lemma 4.1), and similarly for Čech complexes (Lemma 2.1).
Our investigations of Hausdorff versus Gromov–Hausdorff distances have touched upon a number of naturally arising open questions which we believe merit further study. We list these questions here:
Question 1.
The constant $\alpha(n,\kappa)$ in Theorem 4 attains the value $1$ when $M$ is the circle, which cannot be improved. Are the constants $\alpha(n,\kappa)$ optimal in some sense for more general manifolds $M$?
Question 2.
In Theorem 2, we have $d_{\mathrm{GH}}(X,S^{1})=d_{\mathrm{H}}(X,S^{1})$ for $X\subseteq S^{1}$ with $d_{\mathrm{H}}(X,S^{1})<\frac{\pi}{6}$. Is $\frac{\pi}{6}$ optimal in order to achieve equality, i.e., can we construct an example subset $X$ with density slightly below $\frac{\pi}{6}$ and $d_{\mathrm{GH}}(X,S^{1})<d_{\mathrm{H}}(X,S^{1})$? More generally, what happens to the ratio
$\inf\left\{\frac{d_{\mathrm{GH}}(X,S^{1})}{d_{\mathrm{H}}(X,S^{1})}~{}:~{}X\subseteq S^{1}\text{ with }d_{\mathrm{H}}(X,S^{1})\geq\delta\right\}$
as $\delta$ decreases beneath $\frac{\pi}{6}$?
Question 3.
Our results are for a manifold $M$ equipped with the Riemannian metric. If $M$ is instead a manifold embedded in Euclidean space, and equipped with the Euclidean metric, then can one prove analogous results?
Question 4.
What if $M$ is not a manifold, but instead say a graph, or a stratified space, or a length space? When can we prove that for dense enough samplings $X,Y\subseteq M$, the Gromov–Hausdorff distance $d_{\mathrm{GH}}(X,Y)$ is lower bounded in terms of the Hausdorff distance?
Question 5.
Let $M$ be a connected, closed Riemannian manifold. As explained at the end of Section 5, Theorem 1(b) implies that if $X$ and $Y$ are two isometric subsets, and if $Y$ dense enough ($d_{\mathrm{H}}(Y,M)<\frac{1}{4}\rho(M)$), then the Hausdorff distance from $X$ to $M$ and from $Y$ to $M$ cannot differ by more than a factor of two: $d_{\mathrm{H}}(X,M)\leq 2\cdot d_{\mathrm{H}}(Y,M)$. The example in Figure 6 shows that a density assumption like $d_{\mathrm{H}}(Y,M)<\frac{1}{4}\rho(M)$ is needed, since otherwise we could have $d_{\mathrm{H}}(X,M)\approx 4\cdot d_{\mathrm{H}}(Y,M)$. For $X$ and $Y$ two isometric subsets of $M$ with no density assumption, how large can the ratio $\sup\left\{\tfrac{d_{\mathrm{H}}(X,M)}{d_{\mathrm{H}}(Y,M)}~{}|~{}0<d_{\mathrm{H}}(Y,M)\leq d_{\mathrm{H}}(X,M)\right\}$ be?
Question 6.
Are there variants of our results (Theorems 13, and 4) for subsets of a complete manifold without boundary (that could be unbounded)? For example, are there variants of our results for subsets of $\mathbb{R}^{n}$? Does Borel–Moore homology (or alternatively cohomology with compact support), in which every oriented manifold has a fundamental class, play a role?
Question 7.
As a specific example of the question above, are there versions of our results that hold for two periodic point sets in $\mathbb{R}^{n}$, where a periodic point set is defined as the image of a set of points in $[0,2\pi]^{n}$ under the natural action of $(2\pi\mathbb{Z})^{n}$ on $\mathbb{R}^{n}$? One can recast the periodic point set as a point set in the Torus $T^{n}$, and then proceed as in Example 7.5. But it is not clear whether non-periodic correspondences on the point sets on $\mathbb{R}^{n}$ can have smaller distortion than periodic correspondences induced from correspondences on point sets in $T^{n}$.
More generally, let $M$ be a complete manifold without boundary. Suppose $G$ is a group that acts properly and by isometries on $M$, and that $X,Y\subseteq M$ are $G$-invariant subsets. Then the quotient space $M/G$ is a metric space that contains $X/G$ and $Y/G$. What can be said about the relationship between $d_{\mathrm{GH}}(X/G,Y/G)$ and $d_{\mathrm{GH}}(X,Y)$, especially for $X,Y\subseteq M$ sufficiently dense? We note that if $M/G$ is furthermore a compact manifold without boundary, then Theorem 1 applies to give that
$d_{\mathrm{GH}}(X/G,Y/G)$ $\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X/G,M/G)-d_{\mathrm{H}}(Y/G,X/G),\tfrac{1}{6}\rho(M/G)-\tfrac{2}{3}d_{\mathrm{H}}(Y/G,M/G)\right\}$
$=\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\tfrac{1}{6}\rho(M/G)-\tfrac{2}{3}d_{\mathrm{H}}(Y,M)\right\}.$
Question 8.
What if $M$ is not a closed manifold, but instead a compact manifold with boundary $\partial M\neq\emptyset$?
Let $X,Y\subseteq M$ with $\partial M\neq\emptyset$. One thing that could be done in this situation is to define $N$ to be the closed manifold ($\partial N=\emptyset$) obtained by gluing two copies of $M$ together along their common boundary. Now, define $X_{N},Y_{N}\subseteq N$ to be the subsets obtained by duplicating $X$ and $Y$ (although points in the boundary of $M$ will not be duplicated). Suppose furthermore that $N$ is a closed Riemannian manifold with convexity radius $\rho(N)$. Then Theorem 1 applies to give that
$d_{\mathrm{GH}}(X_{N},Y_{N})$ $\geq\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X_{N},N)-d_{\mathrm{H}}(Y_{N},N),\tfrac{1}{6}\rho(N)-\tfrac{2}{3}d_{\mathrm{H}}(Y_{N},N)\right\}$
$=\min\left\{\tfrac{1}{2}d_{\mathrm{H}}(X,M)-d_{\mathrm{H}}(Y,M),\tfrac{1}{6}\rho(N)-\tfrac{2}{3}d_{\mathrm{H}}(Y,M)\right\}.$
What can be said about the relationship between $d_{\mathrm{GH}}(X,Y)$ and $d_{\mathrm{GH}}(X_{N},Y_{N})$, especially for $X,Y\subseteq M$ sufficiently dense?
Question 9.
For $X,Y\subseteq\mathbb{R}^{n}$, there is a nesting $d_{\mathrm{GH}}(X,Y)\leq d_{\mathrm{H,iso}}^{\mathbb{R}^{n}}(X,Y)\leq d_{\mathrm{H}}(X,Y)$, where $d_{\mathrm{H,iso}}^{\mathbb{R}^{n}}$ allows one to align $X$ and $Y$ only via Euclidean isometries; see [43, 44]. Similarly, if $M$ is a manifold with symmetries, and $X,Y\subseteq M$, there is a nesting $d_{\mathrm{GH}}(X,Y)\leq d_{\mathrm{H,iso}}^{M}(X,Y)\leq d_{\mathrm{H}}(X,Y)$, where $d_{\mathrm{H,iso}}^{M}$ allows one to align $X$ and $Y$ only via isometries of the manifold $M$. Can the techniques in our paper be extended to place better bounds on $d_{\mathrm{H,iso}}^{M}$?
Acknowledgements
HA and SM would like to thank the Department of Mathematics at the University of Florida for hosting a research visit. HA was funded in part by the Simons Foundation’s Travel Support for Mathematicians program. FF was funded in part by NSF CAREER Grant DMS 2042428.
Appendix A Proofs of interleaving lemmas
This appendix contains the proofs of Lemmas 4.1 and 4.2. We include these proofs here so that our paper is self-contained, but we also refer the reader to [15] for more details.
Proof of Lemma 4.1.
Let $X$ and $Y$ be metric spaces and let $r>2\cdot d_{\mathrm{GH}}(X,Y)$. We must show that for any $\varepsilon>0$, there exist simplicial maps:
[…]
such that the composition $\overline{h}\circ\overline{g}$ is contiguous to the inclusion $\iota:\mathrm{VR}\left(Y;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(Y;2r+\varepsilon\right)$.
Since $2\cdot d_{\mathrm{GH}}(X,Y)<r$, there exists a correspondence $C\subseteq X\times Y$ with $\mathrm{dis}(C)<r$. We can define (possibly discontinuous) functions $g\colon X\to Y$ and $h\colon Y\to X$ such that $(x,g(x))\in C$ for any $x\in X$ and $(h(y),y)\in C$ for any $y\in Y$.
We first show that the map $h$ extends to a simplicial map $\overline{h}\colon\mathrm{VR}\left(Y;\varepsilon\right)\to\mathrm{VR}\left(X;r+\varepsilon\right)$. Take a simplex $\sigma=[y_{0},\ldots,y_{k}]\in\mathrm{VR}\left(Y;\varepsilon\right)$. By the definition of Vietoris–Rips complex, the diameter of $\sigma$ is smaller than than $\varepsilon$, i.e., $d_{Y}(y_{i},y_{j})<\varepsilon$ for $0\leq i,j\leq k$. Since $(h(y_{i}),y_{i}),(h(y_{j}),y_{j})\in C$, by the definition of distortion we have
$d_{X}(h(x_{i}),h(x_{j}))\leq d_{Y}(y_{i},y_{j})+r<r+\varepsilon.$
So, the diameter of $\overline{h}(\sigma)=[h(y_{0}),\ldots,h(y_{k})]$ is smaller than $r+\varepsilon$, implying that $\overline{h}(\sigma)\in\mathrm{VR}\left(X;r+\varepsilon\right)$, and so $\overline{h}$ is a simplicial map. The proof that $g$ extends to a simplicial map $\overline{g}$ is analogous.
To show that the simplicial maps $\overline{g}\circ\overline{h}$ and $\iota$ are contiguous, we take an arbitrary $k$-simplex $\sigma=[y_{0},\ldots,y_{k}]\in\mathrm{VR}\left(Y;\varepsilon\right)$, meaning that $d_{Y}(y_{i},y_{j})<\varepsilon$ for all $i,j$. Note that $\iota(\sigma)\in\mathrm{VR}\left(Y;2r+\varepsilon\right)$ and $(\overline{g}\circ\overline{h})(\sigma)\in\mathrm{VR}\left(X;2r+\varepsilon\right)$. For any $i,j$, we also have
[…]
$<2r+\varepsilon.$
As a result, the simplex
$\iota(\sigma)\cup(\overline{g}\circ\overline{h})(\sigma)=[y_{0},\ldots,y_{k},g\circ h(y_{0}),\ldots,g\circ h(y_{k})]$
is a simplex in $\mathrm{VR}\left(Y;2r+\varepsilon\right)$. Therefore, the maps are contiguous. ∎
Proof of Lemma 4.2.
Let $Z$ be a metric space, and let $Y\subseteq Z$ satisfy $r>2\cdot d_{\mathrm{H}}(Y,Z)$. We must show that for any $\varepsilon>0$, there exist simplicial maps:
[…]
such that the composition $\iota\circ\overline{f}$ is contiguous to the inclusion $\mathrm{VR}\left(Z;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(Z;r+\varepsilon\right)$. The map $\iota$ is the natural inclusion.
Since $d_{\mathrm{H}}(Y,Z)<\frac{r}{2}$, we can define a (possibly discontinuous) function $f\colon Z\to Y$ such that $d_{Z}(z,f(z))<\frac{r}{2}$ for any $z\in Z$.
We first show that the map $f$ extends to a simplicial map $\overline{f}\colon\mathrm{VR}\left(Z;\varepsilon\right)\to\mathrm{VR}\left(Y;r+\varepsilon\right)$. For $\sigma=[z_{0},\ldots,z_{k}]\in\mathrm{VR}\left(Z;\varepsilon\right)$, we have
$d_{Z}(f(z_{i}),f(z_{j}))\leq d_{Z}(f(z_{i}),z_{i})+d_{Z}(z_{i},z_{j})+d_{Z}(z_{j},f(z_{j}))<\tfrac{r}{2}+\varepsilon+\tfrac{r}{2}=r+\varepsilon.$
So, the diameter of $\overline{f}(\sigma)=[f(z_{0}),f(z_{1}),\ldots,f(z_{k})]$ is smaller than $r+\varepsilon$, giving $\overline{f}(\sigma)\in\mathrm{VR}\left(Y;r+\varepsilon\right)$. Therefore, $\overline{f}$ is simplicial map.
In order to show that the simplicial maps $(\iota\circ\overline{f})$ and $\iota^{\prime}\colon\mathrm{VR}\left(Z;\varepsilon\right)\hookrightarrow\mathrm{VR}\left(Z;r+\varepsilon\right)$ are contiguous, we take an arbitrary simplex $\sigma=[z_{0},\ldots,z_{k}]\in\mathrm{VR}\left(Z;\varepsilon\right)$, i.e., $d_{Z}(z_{i},z_{j})<\varepsilon$ for all $i,j$. Since $\overline{f}$ is a simplicial map, we have $d_{Z}(f(z_{i}),f(z_{j}))<r+\varepsilon$. We also note that for any $i,j$,
$d_{Z}(z_{i},\iota(f(z_{j})))=d_{Z}(z_{i},f(z_{j}))\leq d_{Z}(z_{i},z_{j})+d_{Z}(z_{j},f(z_{j}))<\varepsilon+\tfrac{r}{2}<\varepsilon+r.$
As a result, the simplex $\iota^{\prime}(\sigma)\cup(\iota\circ\overline{f})(\sigma)=[z_{0},z_{1},\ldots,z_{k},f(z_{0}),f(z_{1}),\ldots,f(z_{k})]$ is in $\mathrm{VR}\left(Z;\varepsilon+2r\right)$. Therefore, the maps are contiguous. ∎
Appendix B Additional proofs
Proof of Lemma 7.3.
Let $A\subseteq M$ be compact, with $\mathrm{diam}(A)<\tau$, where $\tau=\rho(M)$ for $\kappa\leq 0$ and $\tau=\min\left\{\rho(M),\tfrac{\pi}{2\sqrt{\kappa+1}}\right\}$ for $\kappa>0$. We must show that the circumcenter $c(A)$ exists in $M$, and that the diameter and circumradius satisfy $\mathrm{diam}(A)\geq 2\alpha(n,\kappa)\cdot R(A)$.
Since $\tfrac{\pi}{2\sqrt{\kappa+1}}<\tfrac{2\pi}{3\sqrt{\kappa}}$, the existence of a circumcenter $c(A)$ follows immediately from Theorem 7.1. For the diameter bound, we now consider the following two cases:
Case I ($\boldsymbol{\kappa\leq 0}$). This follows from Theorem 7.1, since $\kappa$ is an upper bound on curvature.
Case II ($\boldsymbol{\kappa>0}$). From the definition of circumradius, we observe that $R(A)\leq\mathrm{diam}(A)$.
From Theorem 7.1, we have
$\mathrm{diam}(A)$ $\geq\tfrac{2}{\sqrt{\kappa}}\sin^{-1}\left(\sqrt{\tfrac{n+1}{2n}}\sin\left(\sqrt{\kappa}\;R(A)\right)\right)$
[…]
$=2\sqrt{\tfrac{n+1}{2n}}\ \frac{\sin\left(\sqrt{\kappa}\;R(A)\right)}{\sqrt{\kappa}R(A)}R(A).$
Since $\frac{\sin{x}}{x}$ is a decreasing function for $x\in[0,\pi/2]$, the minimum occurs at the right endpoint of the domain of $x$. Since $R(A)\in\left[0,\tfrac{\pi}{2\sqrt{\kappa+1}}\right]$, we get
$\frac{\sin\left(\sqrt{\kappa}\;R(A)\right)}{\sqrt{\kappa}R(A)}\geq\frac{\sin\left(\pi\sqrt{\kappa}/2\sqrt{\kappa+1}\right)}{\pi\sqrt{\kappa}/2\sqrt{\kappa+1}}=\sin\left(\tfrac{\pi}{2}\sqrt{\tfrac{\kappa}{\kappa+1}}\right)/\left(\tfrac{\pi}{2}\sqrt{\tfrac{\kappa}{\kappa+1}}\right).$
Therefore,
$\mathrm{diam}(A)\geq 2\sqrt{\tfrac{n+1}{2n}}\cdot\left(\sin\left(\tfrac{\pi}{2}\sqrt{\tfrac{\kappa}{\kappa+1}}\right)/\left(\tfrac{\pi}{2}\sqrt{\tfrac{\kappa}{\kappa+1}}\right)\right)\cdot R(A)=2\alpha(n,\kappa)\cdot R(A).$